# Optical turbulence profiling at the Table Mountain Facility with the Laser Communication Relay Demonstration GEO downlink

### v1

@@
journal: opticajournalarticletype: Research Article{abstract*}
We present the first time the profile of atmospheric optical turbulence has been measured using the transmitted beam from a satellite laser communication terminal. A Ring Image Next Generation Scintillation Sensor (RINGSS) instrument for turbulence profiling, as described in Tokovinin (MNRAS, 502.1, 2021), was deployed at the NASA/Jet Propulsion Laboratory’s Table Mountain Facility (TMF) in California [1]. The optical turbulence profile was measured with the downlink optical beam from the Laser Communication Relay Demonstration (LCRD) Geostationary satellite. LCRD conducts links with the Optical Communication Telescope Laboratory ground station and the RINGSS instrument was co-located at TMF to conduct measurements. Turbulence profiles were measured at day and night and atmospheric coherence lengths were compared with other turbulence monitors such as a solar scintillometer and Polaris monitor. RINGSS sensitivity to boundary layer turbulence, a feature not provided by many profilers, is also shown to agree well with a boundary layer scintillometer at TMF. Diurnal evolution of optical turbulence and measured profiles are presented. The robust correlation of RINGSS with other turbulence monitors demonstrates the concept of free-space optical communications turbulence profiling, which could be adopted as a way to support optical ground stations in a future Geostationary feeder link network. These results also provide further evidence that RINGSS, a relatively new instrument concept, is effective even in strong daytime turbulence and with reasonable ground layer sensitivity.
@@ 1 Introduction
Turbulence degrades free-space optical communication (FSOC) links, particularly for high data rates and coherent links where technologies like adaptive optics (AO) are often employed to compensate [2, 3]. Monitoring atmospheric turbulence, particularly with the detailed measurements of a turbulence profiler, can support the operation and design of AO systems or the forecasting of turbulence conditions, both important capabilities for future FSOC networks [4, 5]. Geostationary orbit (GEO) feeder links have been proposed as way to implement global FSOC networks and numerous countries are progressing this capability [6, 7, 8].
@@
The Laser Communications Relay Demonstration (LCRD) is a payload that was launched to GEO in 2021 onboard the STPSat-6 satellite to demonstrate a FSOC relay, conducting simultaneous bi-directional links with the Optical Communications Telescope Laboratory (OCTL) at Table Mountain Facility (TMF) and another optical ground station at Haleakala, Hawaii [9].
@@
We demonstrate the concept of FSOC turbulence profiling by using the downlink signal from LCRD near $1550$ nm as opposed to celestial objects typically used for turbulence monitors. The narrow linewidth from the satellite downlink allows narrowband filtering for 24 hr measurements and high signal-to-noise down to very shallow solar-separation angles while the stability of GEO allows for long observation periods with minimal pointing and tracking. Furthermore, FSOC turbulence profiling inherently observes through a slant path parallel to the OGS conducting the link, which is beneficial when interpreting these results for OGS operation.
@@
We have deployed a Ring Image Next Generation Scintillation Sensor (RINGSS) at TMF to conduct the turbulence profiling with the LCRD downlink beam. RINGSS was first outlined in [1] as an evolution of the well-established MASS profiler [10]. RINGSS is simple and low-cost in comparison with most previous profiling instruments and its applicability to FSOC is highlighted in [11]. RINGSS can be implemented with a telescope as small as 5" using commercial off-the-shelf components, unlike the more well-known MASS or SCIDAR [12]. RINGSS produces thin ring images when defocused which can be used to estimate the vertical profile of turbulence if images are captured at high enough speeds to resolve turbulent motion as azimuthal variation around the ring, explained in subsection 2.1.
@@ 2 RINGSS at Table Mountain Facility
A narrow linewidth filter at $1550$ nm is placed in the lens barrel before the reimaging optics, ensuring high signal strength at all times of day and even with shallow solar separation. The bright downlink beam from LCRD allows 200 µs exposure times, though the frame rate is limited to 700 Hz. This very short exposure time reduces the blurring of turbulent motion to improve accuracy, particularly in the presence strong turbulence. We house the instrument in a sliding-roof enclosure at TMF close to OCTL, shown in Figure 1. This proximity ensures both systems are observing through a similar atmospheric channel. As LCRD is in a equatorial geostationary orbit, both RINGSS and OCTL statically point at approximately ${}^{\circ}$ in elevation during links with minimal tracking and pointing required. However, the optics of this system induces strong coma that requires tight centering within ±$\pm 10$ arcseconds of the optical axis.
Although the LCRD downlink is bright enough to easily observe the ring with a smaller aperture, constraints of sizing the ring image with small format and large pixel size (12.5 µm) InGaAs arrays makes a 25cm telescope worthwhile. The RINGSS reimaging optics resize the pixel scale to achieve an optimal conjugation height as described in [1]. The optics are also designed to achieve the $1:10$ ratio of spherical aberration to defocus necessary for a conic wavefront, i.e. producing a close to “diffraction-limited” ring width of 2 pixels (Figure 2). Ring width ($R_{\text{width}}$) and radius ($R_{\text{radius}}$) are critical parameters for the instrument, the latter is kept to $10\pm 1$ pixels while $R_{\text{width}}$ is generally maintained as $<3$ pixels.
Figure 2 shows an idealised ring image from our specifications compared with a sample of data from LCRD. In stronger turbulence conditions $R_{\text{width}}$ increases from the desired width, partially due to the blurring of radial motion. Image quality parameters, e.g. $R_{\text{width}}$, $R_{\text{radius}}$, and coma, are used in the software as inputs for auto-focusing, tracking, or rejecting erroneous data.
@@ 2.1 RINGSS background
RINGSS captures defocused ring images of the LCRD downlink at high speed ($700$ Hz), measuring scintillation as azimuthal variations of intensity around the ring over time. Scintillation around the ring image is decomposed into angular frequencies which when combined with theoretical relations for how different frequencies propagate, allows for the optical turbulence profile to be estimated. The frequency spectrum of scintillation relates to the perturbed wavefront having travelled through the atmospheric column because a defocused image approximates the pupil. A brief explanation of the operational principle is provided below but see [1] for a comprehensive explanation. The angular power spectrum (APS), as measured from scintillation around the ring, is retrieved with annular masks corresponding to discrete angular frequencies. For a given angular frequency, $m$, masks applied to the high speed series of images produce a term, $\langle|a_{m}|^{2}\rangle$, to recover the power of that frequency [1]. The value $\langle|a_{m}|^{2}\rangle$ can be thought of as the azimuthal temporal variance through a frequency filter, hence the set of $\langle|a_{m}|^{2}\rangle$ over a range of $m$ represents the APS. The APS is measured in real-time every few seconds from sets of 2000 images.
With the APS measured, RINGSS then relies on a principle similar to MASS and other scintillation-based turbulence profilers, weighting functions can be formulated to describe how spatial frequencies propagate and instrument response to those optical aberrations. Optical turbulence at any altitude is assumed to have a Kolmogorov spectrum with a strength given by the coefficient of the refractive index structure function, $C_{n}^{2}$, in units $m^{-2/3}$. The general principle for weighting functions is that higher frequencies dissipate faster than lower frequencies, such that the shape of the APS can indicate the height of turbulence [10, 13, 14, 15, 1]. The process of computing weighting functions is best explained by Tokovinin in [1], however an important consideration here is that the LCRD downlink is monochromatic, simplifying the process by ignoring polychromatic corrections implemented for prior instruments [16, 14]. Weighting functions are also dependent on conjugation height, itself a function of defocus or the ring radius, e.g. $R_{\text{radius}}=10\pm 1$ pixels corresponds to conjugation heights of $\approx 780\pm 100$, so drifts in defocus can lead to substantial error particularly for turbulence near the surface. We minimise this by building a look-up table of $W_{m}(z)$ for different conjugation heights, which is matched to $R_{\text{radius}}$ measured in real-time. Figure 3 shows a number of weighting functions for $m=2,4$ and $8$ at the nominal defocus ($R_{\text{radius}}=10$ pixels) with inset masks to highlight the amplitude and phase components of filters propagated in the process of computing $W_{m}(z)$.
@@
Figure 3: Weighting functions, $W_{m}(z)$ from 00 to $16$ km for $m=2,4,8$ with displays of the phase ($F_{\varphi}$) and amplitude ($F_{\chi}$) frequency filters that are used for building the instrument response function and computing weighting functions. The y-axis is representative of a contribution normalised over all frequencies, e.g. increasing values of $m=2$ with distance
@@
Once we derive the set of $J_{i}$, we compute a number of integrated parameters from the turbulence profile. The atmospheric coherence length or Fried parameter ($r_{0}$) describes the general turbulence conditions and is used for comparison with other turbulence monitors. The method of computing $r_{0}$ from $J_{i}$ is given in Equation 3 and referred to as $r_{0,\text{scint}}$ to differentiate from the separate method of estimating $r_{0}$ described in subsection 2.2. The value $\cos{\left(\zeta\right)}\sum_{i=0}^{N-1}J_{i}$ in Equation 3 represents the airmass-corrected integral of the turbulence profile with zenith angle, $\zeta$, which is simply the site latitude when observing an equatorial orbit. The $\cos{(\zeta)}$ term takes the turbulence measured through the slant path and scales it to represent turbulence at zenith. All $r_{0}$ values throughout this paper are reported for $\lambda=500$ nm and airmass-corrected to zenith in keeping with the convention for generalising turbulence conditions. Other parameters such as the atmospheric coherence time and isoplanatic angle can also be recovered from the turbulence profile and used for additional purposes such as informing system design, e.g. AO speed requirements or point-ahead compensation as discussed in section 6.
@@ 3 Calibration
A preliminary dataset with RINGSS at TMF of approximately 5 days is used in this paper over a number of LCRD links that were conducted sporadically at daytime and nighttime. In this section we explain how a number of parameters such as $R_{\text{width}}$, $R_{\text{radius}}$, and coma can be used to assess data reliability or tune operation. While inputs for focus and pointing control loops are important, maintaining data integrity throughout the processing pipeline is also critical as sudden drops in quality can be common, e.g. from very high winds near the telescope.
@@
Once data cubes are captured with RINGSS and the APS is measured, images outside the specification of $119<R_{\text{radius}}<$ 11 and $R_{\text{width}}<3.5$, are flagged as low quality and ignored when computing profiles for real-time display of the turbulent conditions (but are saved nonetheless for analysis of instrument performance such as presented here). Once profiles are computed we can estimate both $r_{0,\text{scint}}$ and $r_{0,\text{sector}}$, and use the ratio $r_{0,\text{scint}}/r_{0,\text{sector}}$ for a prima facie indication of accuracy. The radial motion method, $r_{0,\text{sector}}$, is partially independent from the measured optical turbulence profile used for estimating $r_{0,\text{scint}}$. This comparison is similar to the benefit offered from the common implementation of a combined MASS-DIMM instrument [13]. This ratio and the norm of the residuals from solving Equation 2, $\sigma_{NNLS}$, can then be used to indicate low quality data. We also compute profiles from approximately 2 minute rolling averages of the APS, which helps to constrain the solution and ensure all layers return a non-zero estimate of $J_{i}$ [14].
@@
Figure 4: Corner plot displaying the correlation between a number of important parameters for data and image quality, measured from preliminary observations at TMF. Diagonal entries show the histogram of values corresponding to each columns parameter. Dashed red line is the line of best fit from a linear regression between parameters and the $R$ values are the Pearson correlation coefficients between parameters. $r_{0}$ is in centimetres, scaled to $500$ nm, as is done throughout this study.
@@
Detecting low-quality data can be difficult if image quality typically degrades in strong turbulence, requiring careful calibration for any relations that might exist. We produce a corner plot, Figure 4, to show correlations between parameters of interest. Figure 4 shows that $r_{0,\text{scint}}/r_{0,\text{sector}}$ is largely independent of turbulence strength but a loose correlation exists for other parameters with increasing turbulence (decreasing $r_{0}$). $\sigma_{NNLS}$ is correlated with image specification parameters in Figure 4, and outliers of $\sigma_{NNLS}$ from the weak relation with $r_{0}$ were often found to correspond with poor image quality, e.g. weak signal from thin cloud or wind shake causing the spot to temporarily leave the region-of-interest. We aim to find more robust correlations between all parameters shown in Figure 4 to optimise performance and maintain data integrity.

### v2

@@
journal: opticajournalarticletype: Research Article
{abstract*}
We report the first measurement of the atmospheric optical turbulence profile using the transmitted beam from a satellite laser communication terminal. A Ring Image Next Generation Scintillation Sensor (RINGSS) instrument for turbulence profiling, as described in Tokovinin (MNRAS, 502.1, 2021, 747-808), was deployed at the NASA/Jet Propulsion Laboratory’s Table Mountain Facility (TMF) in California [1]. The optical turbulence profile was measured with the downlink optical beam from the Laser Communication Relay Demonstration (LCRD) Geostationary satellite. LCRD conducts links with the Optical Communication Telescope Laboratory ground station and the RINGSS instrument was co-located at TMF to conduct measurements. Turbulence profiles were measured at day and night and atmospheric coherence lengths were compared with other turbulence monitors such as a solar scintillometer and Polaris motion monitor. RINGSS sensitivity to boundary layer turbulence, a feature not provided by many profilers, is also shown to agree with a boundary layer scintillometer at TMF ($R=0.85$). Diurnal evolution of optical turbulence and measured profiles are presented. The correlation of RINGSS with other turbulence monitors ($R=0.75-0.86$) demonstrates the concept of free-space optical communications turbulence profiling, which could be adopted as a way to support optical ground stations in a future Geostationary feeder link network. These results also provide further evidence that RINGSS, a relatively new instrument concept, correlates well with other instruments in daytime and nighttime turbulence.
@@ 1 Introduction
The prescence of optical turbulence degrades free-space optical communication (FSOC) links, particularly for high data rates and coherent links where technologies like adaptive optics (AO) are often employed to compensate [2, 3]. Measurements from monitoring atmospheric turbulence have many applications to FSOC, such as supporting the operation and design of AO systems, forecasting of turbulence conditions, or refining link budgets [4, 5]. Improving AO systems will lead to more reliable links by stabilising and boosting received power. High data rate modems require the coupling of the downlink signal into a single mode fiber and are therefore especially sensitive to the received point spread function (PSF). In these configurations, the use of AO brings the downlink PSF closer to the diffraction limit and improves the coupling efficiency [2, 3]. Turbulence forecasting can also be used with a network of optical ground stations (OGS) to help plan space-to-ground links based on future conditions such as received power. Although more complex, resolving the vertical profile of atmospheric turbulence is a superior tool to measuring integrated turbulence for all these applications. The turbulence profile contains more information which can be used to derive unique parameters such as the isoplanatic angle and coherence time, both of which are valuable for designing uplink and adaptive optics systems.
@@
Geostationary orbit (GEO) feeder links have been proposed as way to implement global FSOC networks and numerous countries are progressing this capability [6, 7, 8]. The Laser Communications Relay Demonstration (LCRD) is a payload that was launched to GEO in 2021 onboard the STPSat-6 satellite to demonstrate a FSOC relay, conducting simultaneous bi-directional links with the Optical Communications Telescope Laboratory (OCTL) at Table Mountain Facility (TMF) and another OGS at Haleakala, Hawaii [9].
@@
We demonstrate the concept of FSOC turbulence profiling by using the downlink signal from LCRD near $1550$ nm as opposed to celestial objects typically used for turbulence monitors. The narrow linewidth from the satellite downlink allows narrowband filtering for 24 hr measurements and high signal-to-noise down to very shallow solar-separation angles while the stability of GEO allows for long observation periods with minimal pointing and tracking. Furthermore, FSOC turbulence profiling inherently observes through a slant path parallel to the OGS conducting the link, which is beneficial when interpreting these results for OGS operation. Real-time information about the atmospheric channel could be used for turbulence forecasting and to optimise AO performance. A space terminal could also adapt bit rate depending on the received power from ground terminal forecasts.
@@
We have deployed a Ring Image Next Generation Scintillation Sensor (RINGSS) at TMF to conduct the turbulence profiling with the LCRD downlink beam. RINGSS, first outlined in [1] as an evolution of the well established MASS profiler, operates by capturing defocused ring images at high speed. Scintillation is imaged as azimuthal variation around the defocused ring image and the process of estimating a turbulence profile from this is explained in subsection 2.1. Simpler methods exist for measuring integrated turbulence, such as solar scintillometers, DIMMs, or motion monitors (introduced in section 4), but do not provide the additional capability provided from resolving the vertical profile of turbulence. RINGSS is novel as a profiler that can be implemented with a telescope as small as 12 cm using commercial off-the-shelf components, unlike the more well-known MASS which requires many custom components, or SCIDAR which requires a $\gtrapprox 1$ m telescope [10, 11]. A number of other instruments, namely the SHIMM and FASS, have also been developed recently and show promise for turbulence profiling with reduced complexity [12, 13].
@@ 2 RINGSS at Table Mountain Facility
A narrow linewidth filter at $1550$ nm is placed in the lens barrel before the reimaging optics, ensuring high signal strength at all times of day and even with shallow solar separation. The bright downlink beam from LCRD allows 200 µs exposure times, though the frame rate is limited to 700 Hz. This very short exposure time reduces the blurring of turbulent motion to improve accuracy, particularly in the presence strong turbulence. We house the instrument in a sliding-roof enclosure at TMF close to OCTL, shown in Figure 1. This proximity ensures both systems are observing through a similar atmospheric channel. As LCRD is in a equatorial geostationary orbit, both RINGSS and OCTL statically point at approximately 50 in elevation during links with minimal tracking and pointing required. However, the optics of this system induces strong coma that requires tight centering within ±$\pm 10$ arcseconds of the optical axis.
Although the LCRD downlink is bright enough to easily observe the ring with a smaller aperture, constraints of sizing the ring image with small format and large pixel size (12.5 µm) InGaAs arrays makes a 25cm telescope worthwhile. The RINGSS reimaging optics resize the pixel scale to achieve an optimal conjugation height as described in [1]. The optics are also designed to achieve the $1:10$ ratio of spherical aberration to defocus necessary for a conic wavefront, i.e. producing a close to “diffraction-limited” ring width of 2 pixels (Figure 2). Ring width ($R_{\text{width}}$) and radius ($R_{\text{radius}}$) are critical parameters for the instrument, the latter is kept to $10\pm 1$ pixels while $R_{\text{width}}$ is generally maintained as $<3$ pixels. Figure 2 shows an idealised ring image from our specifications compared with a sample of data from LCRD. Image quality parameters, e.g. $R_{\text{width}}$, $R_{\text{radius}}$, and coma, are used in the software as inputs for auto-focusing, tracking, or rejecting erroneous data.
@@ 2.1 RINGSS background
RINGSS deployed at TMF captures defocused ring images of the LCRD downlink at high speed ($700$ Hz), measuring scintillation as azimuthal variations of intensity around the ring over time. Scintillation around the ring image is decomposed into angular frequencies which when combined with theoretical relations for how different frequencies propagate, allows for the optical turbulence profile to be estimated. The frequency spectrum of scintillation relates to the perturbed wavefront having travelled through the atmospheric column because a defocused image approximates the pupil. A brief explanation of the operational principle is provided below but see [1] for a comprehensive explanation. The angular power spectrum (APS), as measured from scintillation around the ring, is retrieved with annular masks corresponding to discrete angular frequencies. For a given angular frequency, $m$, masks applied to the high speed series of images produce a term, $\langle|a_{m}|^{2}\rangle$, to recover the power of that frequency [1]. The value $\langle|a_{m}|^{2}\rangle$ can be thought of as the azimuthal temporal variance through a frequency filter, hence the set of $\langle|a_{m}|^{2}\rangle$ over a range of $m$ represents the APS. $200$ µs exposures do not warrant a finite-exposure time correction, as used by [1] for 1 ms exposures, $\langle|a_{m}|^{2}\rangle$ values are adjusted for a detector read noise of $30$ $e^{-}$. The APS is measured in real-time every few seconds from sets of 2000 images.
With the APS measured, RINGSS then relies on a principle similar to MASS and other scintillation-based turbulence profilers, weighting functions can be formulated to describe how spatial frequencies propagate and instrument response to those optical aberrations. Optical turbulence at any altitude is assumed to have a Kolmogorov spectrum with a strength given by the coefficient of the refractive index structure function, $C_{n}^{2}$, in units $m^{-2/3}$. The general principle for weighting functions is that higher frequencies decrease with distance faster than lower frequencies, such that the shape of the APS can indicate the height of turbulence [11, 14, 15, 16, 1]. The process of computing weighting functions is best explained by Tokovinin in [1], however an important consideration here is that the LCRD downlink is monochromatic, simplifying the process by ignoring polychromatic corrections implemented for prior instruments [17, 15]. Weighting functions are also dependent on conjugation height, itself a function of defocus or the ring radius, e.g. $R_{\text{radius}}=10\pm 1$ pixels corresponds to conjugation heights of $\approx-780\pm 100$ m, so drifts in defocus can lead to substantial error particularly for turbulence near the surface. We minimise this by building a look-up table of $W_{m}(z)$ for different conjugation heights, which is matched to $R_{\text{radius}}$ measured in real-time. Figure 3 shows weighting functions for $m=2,4$ and $8$ at the nominal defocus ($R_{\text{radius}}=10$ pixels).
@@
Figure 3: Weighting functions, $W_{m}(z)$, from 00 to $16$ km for $m=2,4,8$.
@@
Once we derive the set of $J_{i}$, we compute a number of integrated parameters from the turbulence profile. The atmospheric coherence length or Fried parameter ($r_{0}$) describes the general turbulence conditions and is used for comparison with other turbulence monitors. The method of computing $r_{0}$ from $J_{i}$ is given in Equation 3 and referred to as $r_{0,\text{scint}}$ to differentiate from the separate method of estimating $r_{0}$ described in subsection 2.2. The value $\cos{\left(\zeta\right)}\sum_{i=0}^{N-1}J_{i}$ in Equation 3 represents the airmass-corrected integral of the turbulence profile with zenith angle, $\zeta$, which is approximately $40^{\circ}$ for TMF observing LCRD in Geostationary orbit. The $\cos{(\zeta)}$ term takes the turbulence measured through the slant path and scales it to represent turbulence at zenith. All $r_{0}$ values throughout this paper are reported for $\lambda=500$ nm and airmass-corrected to zenith in keeping with the convention for generalising turbulence conditions. Other parameters such as the atmospheric coherence time and isoplanatic angle can also be recovered from the turbulence profile and used for additional purposes such as informing system design, e.g. AO speed requirements or point-ahead compensation as discussed in section 6.
@@ 3 Calibration
A preliminary dataset with RINGSS at TMF of approximately 5 days throughout May and June of 2023 is used in this paper over a number of LCRD links that were conducted sporadically at daytime and nighttime. These observations correspond to approximately twelve hours of turbulence profile data. In this section we explain how a number of parameters such as $R_{\text{width}}$, $R_{\text{radius}}$, and coma can be used to assess data reliability or tune operation. While inputs for focus and pointing control loops are important, maintaining data integrity throughout the processing pipeline is also critical as sudden drops in quality can be common, e.g. from very high winds near the telescope.
@@
Once data cubes are captured with RINGSS and the APS is measured, images outside the specification of $119<R_{\text{radius}}<$ 11 and $R_{\text{width}}<3.5$, are flagged as low quality and ignored when computing profiles for real-time display of the turbulent conditions (but are saved nonetheless for analysis of instrument performance such as presented here). Once profiles are computed we can estimate both $r_{0,\text{scint}}$ and $r_{0,\text{sector}}$, and use the ratio $r_{0,\text{scint}}/r_{0,\text{sector}}$ for a prima facie indication of accuracy. The radial motion method, $r_{0,\text{sector}}$, is partially independent from the measured optical turbulence profile used for estimating $r_{0,\text{scint}}$. This comparison is similar to the benefit offered from the common implementation of a combined MASS-DIMM instrument [14]. The norm of the residuals from the non-negative least squares solution to Equation 2, $\sigma_{NNLS}$, can also be used to indicate low quality data. We also compute profiles from approximately 2 minute rolling averages of the APS, which helps to constrain the solution and minimise zero estimates of $J_{i}$ [15].
@@
Figure 4: Corner plot displaying the correlation between a number of important parameters for data and image quality, measured from preliminary observations at TMF. Diagonal entries show the histogram of values corresponding to each columns parameter. Dashed red line is the line of best fit from a linear regression between parameters and the $R$ values are the Pearson correlation coefficients between parameters. $r_{0}$ is in centimetres, scaled to $500$ nm and zenith, as is done throughout this study.
@@
Detecting low-quality data can be difficult if image quality typically degrades in strong turbulence, requiring careful calibration for any relations that might exist. We produce a corner plot, Figure 4, to show correlations between parameters of interest. Figure 4 shows that $r_{0,\text{scint}}/r_{0,\text{sector}}$ is largely independent of turbulence strength but a loose correlation exists for other parameters with increasing turbulence (decreasing $r_{0}$). $\sigma_{NNLS}$ is correlated with image specification parameters in Figure 4, and outliers of $\sigma_{NNLS}$ from the weak relation with $r_{0}$ were often found to correspond with poor image quality, e.g. weak signal from thin cloud or wind shake causing the spot to temporarily leave the region-of-interest. We also observe $R_{\text{width}}$ increases in stronger turbulence, possibly due to the blurring of rapid changes in $R_{\text{radius}}$ within exposures, but may also be caused by high wind speed. We aim to find more robust correlations between all parameters shown in Figure 4 to optimise performance and maintain data integrity.